OpenStudy (anonymous):

Find the break-even point by solving each system using substitution. Show all work. 20. Cost: y = x + 100 Revenue: y = 3x 3x=x+100 2x=100 x=50 Break-even Point:

OpenStudy (anonymous):

@jdoe0001 @jim_thompson5910 any chance someone can show me how to find the break even point?@wolf1728

OpenStudy (wolf1728):

Any more information with that problem?

OpenStudy (wolf1728):

I guess you are asking to solve these 2 equations? y = x + 100 y = 3x

OpenStudy (anonymous):

um im not sure it says to find the break even point, when I solved those equations I got x=50 and y =150 but which is the breakeven point?

OpenStudy (wolf1728):

So they just give you the equations and you have to figure what is calculating what?

OpenStudy (anonymous):

?

OpenStudy (wolf1728):

Okay, I'm guessing you have to determine at what point this becomes profitable. So we have the cost expressed as y = 50 + 100 revenue is calculated at: y = 3 * x I'm guessing that it costs 150 to manufacture something and you have to sell it at 3x. The breakeven point must be where the items sold yields enough revenue to match all the money spent for cost. And there's no more information on this?

OpenStudy (anonymous):

no but thatll do ty

OpenStudy (wolf1728):

well thank you :-)